# BranchGRPO: Stable and Efficient GRPO with Structured Branching in Diffusion Models

### v1

@@ Abstract
Recent advancements in aligning image and video generative models via GRPO have achieved remarkable gains in enhancing human preference alignment. However, these methods still face high computational costs from on-policy rollouts and excessive SDE sampling steps, as well as training instability due to sparse rewards. In this paper, we propose BranchGRPO, a novel method that introduces branch sampling policy that updates SDE sampling process. By sharing computation across common prefixes and pruning low-reward paths and redundant depths, BranchGRPO substantially lowers the per-update compute cost while maintaining or improving exploration diversity. This work makes three main contributions: (1) a branch sampling scheme that reduces the cost of rollouts and training; (2) a tree-based advantage estimator that incorporates dense process-level rewards; and (3) pruning strategies that exploit path and depth redundancy to accelerate convergence and boost performance. Experiments on image and video preference alignment show that BranchGRPO improves alignment scores by 16% over strong baselines, while cutting training time by 50%.
@@ 1 Introduction
Diffusion and flow-matching models have advanced image and video generation with high fidelity, diversity, and controllability (Ho et al., 2020a; Rombach et al., 2022b; Lipman et al., 2022). However, large-scale pretraining alone cannot ensure alignment with human intent, as outputs often miss aesthetic, semantic, or temporal expectations. Reinforcement learning from human feedback (RLHF) addresses this gap by directly adapting models toward human-preferred outcomes (Ouyang et al., 2022).
Within RLHF, Group Relative Policy Optimization (GRPO) has shown strong stability and scalability across text-to-image and text-to-video tasks (Shao et al., 2024; Xue et al., 2025). However, when applied to diffusion and flow-matching models, current GRPO variants still face two fundamental bottlenecks: (1) Inefficiency. Standard GRPO adopts a sequential rollout design, where each trajectory must be independently sampled under both the old and new policies. This incurs $O(N\cdot T)$ complexity with denoising steps $T$ and group size $N$, leading to significant computational redundancy and limiting scalability in large-scale image and video generation tasks. (2) Sparse rewards. Most methods compute only a single reward from the final denoised image or video and propagate it uniformly across all denoising steps. This trajectory-level supervision neglects the informative signals contained in intermediate states, resulting in unstable optimization, inaccurate credit assignment, and high-variance gradients. This raises a central challenge: how can we accurately attribute sparse outcome rewards to the specific denoising steps that truly contribute to high-quality generations?
To overcome these limitations, we propose BranchGRPO, a tree-structured policy optimization framework for diffusion and flow models. BranchGRPO replaces inefficient independent sequential rollouts with a branching structure, where scheduled split steps in the denoising process allow each trajectory to stochastically expand into multiple sub-trajectories while reusing shared prefixes. This design amortizes computation across common segments, substantially improves sampling efficiency, and significantly reduces computational cost.
@@
Our contributions can be summarized as follows:
  • We introduce BranchGRPO, a new training framework that replaces independent sequential rollouts with a tree-structured branching mechanism. By inserting branching points during the denoising process, trajectories can expand into multiple sub-trajectories while sharing early prefixes, which substantially reduces redundancy and improves exploration efficiency.
@@
  • We design complementary width- and depth-pruning strategies that reduce backpropagation cost and further improve alignment scores.
@@ 2 Related Work
The past few years have witnessed rapid progress in generative modeling of visual content. Diffusion models Ho et al. (2020b); Rombach et al. (2022a) and, more recently, flow-matching models Lipman et al. (2022); Liu et al. (2023) provide strong unconditional image generators, yet aligning them with nuanced human aesthetic and semantic preferences remains a challenge. Reinforcement learning from human feedback (RLHF) was first explored in natural language processing to fine-tune language models toward user desires Christiano et al. (2017); Ziegler et al. (2019); Stiennon et al. (2020); Ouyang et al. (2022). Inspired by this success, researchers have adapted RLHF concepts to visual generative models by training learned reward functions and backpropagating differentiable rewards through the denoising process Clark et al. (2023); Prabhudesai et al. (2024). Other works build reward models from large user comparison datasets or develop multi-faceted video reward functions Liu et al. (2025a), enabling RL-based post-training that improves realism and text-image alignment.
Parallel to reward-model-based RLHF, policy-gradient methods treat the diffusion sampling procedure as a multi-step decision process. The Denoising Diffusion Policy Optimization (DDPO) framework of Black et al. Black et al. (2023) demonstrated that Proximal Policy Optimization can effectively fine-tune diffusion models on learned rewards beyond prompt engineering. Fan et al. later introduced DPOK Fan et al. (2023), which adds KL-regularized updates to stabilize training on text-to-image tasks. Subsequent works explored self-play fine-tuning Yuan et al. (2024) and aesthetic post-training with step-wise preference optimization Liang et al. (2025). Direct Preference Optimization (DPO) bypasses explicit reward modeling by directly training on comparison data Rafailov et al. (2023), and several works have extended this idea to diffusion: D3PO for diffusion models Yang et al. (2024) and Diffusion-DPO Wallace et al. (2024) achieve comparable or better alignment than reward-model methods while reducing complexity. For video, online preference optimization algorithms such as OnlineVPO Zhang et al. (2024) and related human-feedback-driven methods Lee et al. (2023) are emerging to align video diffusion generators.
Recently, Group Relative Policy Optimization (GRPO) has offered a promising alternative to PPO by leveraging relative rewards across a batch of sampled outputs Guo et al. (2024). GRPO obviates the need for learned critics and yields stable updates in language model post-training. DanceGRPO Xue et al. (2025) extends this approach to both diffusion and flow-matching models, introducing grouped rollouts and converting deterministic flows to equivalent SDEs for exploration. Flow-GRPO Liu et al. (2025b) applies GRPO specifically to flow-matching generators, proposing denoising-step reduction to improve efficiency and demonstrating strong gains on compositionality and text-rendering tasks. MixGRPO Li et al. (2025) further reduces overhead by mixing ODE and SDE steps, though its impact is comparatively modest. Building on these advances, our work continues to explore group-based RL for aligning image and video generators with human preferences.
@@ 3.2 Branch Rollout Algorithm
Given a prompt, BranchGRPO reformulates denoising into a tree-structured process. We align terminology with tree search: (i) depth $T$ denotes denoising steps; (ii) width $w$ counts completed trajectories (leaves); (iii) branching steps $\mathcal{B}$ indicate split steps; (iv) branch correlation $s$ controls inter-branch noise correlation; (v) branching factor $K$ is the number of children per split.
@@
where $\xi_{0},\delta\xi_{b}\sim\mathcal{N}(0,I)$ and $s\!\geq\!0$ adjusts inter-branch correlation. When $s=0$, all branches collapse to a single path, while as $s\!\to\!\infty$, the branches become independent.
Figure 3: Illustration of BranchGRPO training. Left: Branch rollout. A root trajectory expands at split points into multiple children, yielding diverse leaf samples with associated rewards. Middle: Reward fusion. Leaf rewards are propagated upward and aggregated at internal nodes using path-probability weighting. Right: Depth-wise normalization and pruning. Aggregated rewards are standardized within each layer to compute structured advantages, while selected nodes can be pruned (red) to reduce training cost. This process converts sparse terminal rewards into dense, depth-specific signals aligned with the tree topology.
@@ 3.3 Reward Fusion and Depth-wise Advantage Estimation
where $s_{\ell}$ is the path log-probability under the behavior policy and $\beta$ controls the strength of weighting. When $\beta=0$, this reduces to simple averaging; larger $\beta$ emphasizes high-likelihood paths and reduces the influence of noisy, low-probability leaves.
@@ 3.4 Pruning Strategies
While branch rollouts improve efficiency and provide dense process rewards, an excessive number of branches may induce exponential growth in trajectory count, leading to prohibitive backpropagation cost. Recent work (Li et al., 2025) has revealed that standard GRPO training contains substantial redundancy and pruning potential. Motivated by this, we introduce two complementary pruning strategies in the context of BranchGRPO: width pruning, which reduces the number of leaves used for gradient updates, and depth pruning, which skips unnecessary denoising layers.
@@
After computing rewards and advantages for all leaves $\mathcal{L}$, we select a subset $\mathcal{S}\subseteq\mathcal{L}$ for gradient updates:
$\mathcal{S}\;=\;\mathrm{SelectWidth}\!\left(\mathcal{L};\ \texttt{mode}_{w},\ K\right),\qquad|\mathcal{S}|=K\leq|\mathcal{L}|.$ (6)
We consider two simple selection modes:
  • Parent-Top1: keep the better child of each parent in the last branching step. This halves gradient cost while ensuring that every parent is represented, leading to more stable updates but with reduced diversity since only “winners” are kept.
  • Extreme selection: retain the globally best and worst $b$ leaves by score $s(\ell)$. By preserving both strong positive and negative signals, this strategy provides richer gradient feedback and encourages exploration, though at the cost of higher variance in training.
@@
Notably, recent work MixGRPO (Li et al., 2025) achieves strong performance by hybrid ODE–SDE, effectively acting as a form of depth pruning. This motivates our explicit exploration of depth pruning within BranchGRPO.
@@ 4.1 Experiment Setup
We evaluate BranchGRPO on HPDv2.1 (Wu et al., 2023) (103k training and 400 balanced test prompts). The backbone is FLUX.1-Dev (Batifol et al., 2025), with rollouts augmented by branching and process-dense rewards; we also study a pruned variant (BranchGRPO-Pruning). Baselines include DanceGRPO and MixGRPO under identical settings. We report efficiency (NFE, iteration time, GPU-hours) and quality (HPS-v2.1, PickScore (Kirstain et al., 2023), ImageReward (Xu et al., 2023)). Unless otherwise noted, defaults use depth $20$, branch factor $2$ (16 leaves), and correlation $s\in\{0,1,2,4,8\}$, with all reward models and hyperparameters matched across methods. Additional training details are provided in the supplement.
@@ 4.2 Implementation Details
We set the tree depth to $d=20$ and the branch factor to $b=2$, yielding $16$ leaves per rollout before pruning. The branching steps $\mathcal{B}$ uses three presets: Dense $(0,3,6,9)$ which is default setting, Mixed $(0,5,15,20)$, and Sparse $(0,10,13,19)$. The branch correlation $s$ sweeps $s\in\{0,1,2,4,8\}$. Training runs for 300 optimizer steps with gradient accumulation $g=12$ and per-GPU batch size =2=2, on $16\times$ NVIDIA H20z GPUs. Optimization uses AdamW (learning rate $1\times 10^{-5}$, weight decay $1\times 10^{-4}$) with bf16 precision and EMA weights stored on CPU. For a fair comparison, all GRPO-related hyperparameters are kept identical across methods; full hyperparameter details are provided in the supplementary material.
@@ 4.3 Main Results
[…]
The reward curves in Figure 1 further highlight the stability advantage of BranchGRPO. DanceGRPO with reduced timestep fractions (tf=0.6) suffers from large reward fluctuations and unstable convergence, while the full-timestep variant (tf=1.0) converges more smoothly but at prohibitive cost. BranchGRPO achieves both faster reward growth and more stable training trajectories, confirming that branching rollouts provide denser and more reliable credit assignment. Taken together, these results establish BranchGRPO as a more efficient and stable alternative to existing GRPO variants, yielding superior alignment at a substantially reduced training overhead.
@@ 4.4 Ablation Studies
We next vary the positions of split timesteps (Figure 4(b)). Early splits such as $(0,3,6,9)$ promote faster reward increase in the early stage, whereas later splits like $(9,12,15,18)$ delay exploration and yield lower rewards overall. Intermediate schedules (e.g., $(3,6,9,12)$) balance efficiency and reward quality, suggesting that “front-loaded” splits are generally more effective for exploration.
@@
Figure 4: Ablation on branching hyperparameters without any pruning (width or depth). All runs use the same training setup; the default branch correlation is s=4.0s{=}4.0 unless varied in (a). (a) Branch correlation $s$: moderate correlation (s=4.0s{=}4.0) balances exploration and stability and attains the best reward. (b) Branching steps $\mathcal{B}$: front-loaded split schedules (e.g., $(0,3,6,9)$) yield higher reward than late splits (e.g., $(9,12,15,18)$). (c) Branch density: increasing the number of early splits accelerates early reward gains while reaching similar final levels.
@@ 4.5 Video Generation Results
Quantitatively, BranchGRPO achieves faster convergence and higher rewards, measured by the Video-Align motion quality metric (Liu et al., 2025c), confirming the benefit of branching rollouts for video generation.

### v2

@@ Abstract
Recent progress in aligning image and video generative models with Group Relative Policy Optimization (GRPO) has improved human preference alignment, yet existing approaches still suffer from high computational cost due to sequential rollouts and large numbers of SDE sampling steps, as well as training instability caused by sparse rewards. In this paper, we present BranchGRPO, a method that restructures the rollout process into a branching tree, where shared prefixes amortize computation and pruning removes low-value paths and redundant depths. BranchGRPO introduces three contributions: (1) a branch sampling scheme that reduces rollout cost by reusing common segments; (2) a tree-based advantage estimator that converts sparse terminal rewards into dense, step-level signals; and (3) pruning strategies that accelerate convergence while preserving exploration. On HPDv2.1 image alignment, BranchGRPO improves alignment scores by up to 16% over strong baselines, while reducing per-iteration training time by nearly 55%. On WanX-1.3B video generation, it further achieves higher Video-Align scores with sharper and temporally consistent frames compared to DanceGRPO.
@@ 1 Introduction
Diffusion and flow-matching models have advanced image and video generation with high fidelity, diversity, and controllability (Ho et al., 2020; Lipman et al., 2022; Liu et al., 2022). However, large-scale pretraining alone cannot ensure alignment with human intent, as outputs often miss aesthetic, semantic, or temporal expectations. Reinforcement learning from human feedback (RLHF) addresses this gap by directly adapting models toward human-preferred outcomes (Ouyang et al., 2022).
Within RLHF, Group Relative Policy Optimization (GRPO) has shown strong stability and scalability across text-to-image and text-to-video tasks (Liu et al., 2025a; Xue et al., 2025). However, when applied to diffusion and flow-matching models, current GRPO variants still face two fundamental bottlenecks: (1) Inefficiency. Standard GRPO adopts a sequential rollout design, where each trajectory must be independently sampled under both the old and new policies. This incurs $O(N\cdot T)$ complexity with denoising steps $T$ and group size $N$, leading to significant computational redundancy and limiting scalability in large-scale image and video generation tasks. (2) Sparse rewards. Most methods compute only a single reward from the final denoised image or video and propagate it uniformly across all denoising steps. This trajectory-level supervision neglects informative signals contained in intermediate states, resulting in unstable optimization, inaccurate credit assignment, and high-variance gradients. This raises a central challenge: how can we accurately attribute sparse outcome rewards to the specific denoising steps that truly contribute to high-quality generations?
To overcome these limitations, we propose BranchGRPO, a tree-structured policy optimization framework for diffusion and flow models. BranchGRPO replaces inefficient independent sequential rollouts with a branching structure, where scheduled split steps in the denoising process allow each trajectory to stochastically expand into multiple sub-trajectories while reusing shared prefixes. This design amortizes computation across common segments and aligns naturally with the stepwise nature of denoising, substantially improving sampling efficiency while reducing computational cost.
@@
Our contributions are threefold:
  • We introduce BranchGRPO, a tree-structured rollout framework that replaces independent sequential sampling with branching trajectories. By inserting split points during denoising, trajectories expand into multiple sub-trajectories while sharing early prefixes, which reduces redundancy and improves exploration efficiency.
@@
  • We design complementary width- and depth-pruning strategies that lower backpropagation cost and further improve alignment.
  • We validate the effectiveness of BranchGRPO on both text-to-image and image-to-video alignment tasks, demonstrating its general applicability across modalities.
@@ 2 Related Work
Diffusion models (Ho et al., 2020; Rombach et al., 2022) and flow matching models (Lipman et al., 2022; Liu et al., 2022) have become dominant paradigms for visual generation due to their strong theoretical foundations and ability to generate high-quality content efficiently. While pretraining establishes the generative prior, aligning outputs with nuanced human preferences requires reinforcement learning from human feedback (RLHF). In natural language processing, RLHF has proven highly successful for aligning large language models (LLMs) (Ouyang et al., 2022; Christiano et al., 2017), where methods such as PPO and GRPO enable stable preference-driven post-training. These successes have inspired adaptation of RLHF to vision.
In the visual domain, RLHF for diffusion has been developed along two main directions. Reward-model-based approaches such as ImageReward (Xu et al., 2023) backpropagate learned rewards through the denoising process. Direct Preference Optimization (DPO) (Rafailov et al., 2023) has also been extended to diffusion, leading to Diffusion-DPO (Wallace et al., 2024) and Videodpo (Liu et al., 2025c), which achieve competitive alignment without explicit reward modeling. Policy-gradient formulations such as DDPO (black2023ddpo) and DPOK (Fan et al., 2023) further explore online optimization but often face stability challenges. Meanwhile, standardized reward models including HPS-v2.1 (Wu et al., 2023), VideoAlign (Liu et al., 2025b) enable systematic comparison of alignment algorithms on image and video tasks.
More recently, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) has been introduced as a scalable alternative to PPO for preference optimization. DanceGRPO (Xue et al., 2025) pioneers the application of GRPO to visual generation, unifying diffusion and flow models via SDE reformulation and demonstrating stable optimization across text-to-image, text-to-video, and image-to-video tasks. Flow-GRPO (Liu et al., 2025a) extends GRPO to flow matching models by introducing ODE-to-SDE conversion and denoising reduction for improved efficiency. MixGRPO (Li et al., 2025a) further enhances efficiency with a mixed ODE–SDE sliding-window scheme, though it still faces trade-offs between overhead and performance. Our work continues this line by introducing BranchGRPO, which leverages branching rollouts, depth-wise reward fusion, and structured pruning to improve both stability and efficiency; while related to TreePO in LLMs (Li et al., 2025b), our method adapts tree-structured rollouts specifically to diffusion dynamics.
@@ 3.2 Branch Rollout Algorithm
Given a prompt, BranchGRPO reformulates denoising into a tree-structured process. We align terminology with tree search: (i) depth $T$ denotes the number of denoising steps; (ii) width $w$ is the number of completed trajectories (leaves); (iii) branching steps $\mathcal{B}$ indicate split timesteps; (iv) branch correlation $s$ controls the correlation among child noises; (v) branching factor $K$ is the number of children per split.
@@
where $\xi_{0},\delta\xi_{b}\sim\mathcal{N}(0,I)$ and $s\!\geq\!0$ adjusts inter-branch correlation. When $s=0$, all branches collapse to a single path, while as $s\!\to\!\infty$, the branches become independent.Intuitively, smaller $s$ yields highly correlated children (stable but limited exploration), whereas larger $s$ produces more independent trajectories (diverse but noisier).
Figure 3: Left: branching rollouts generate diverse leaves. Middle: leaf rewards are fused upward by path-probability weighting. Right: depth-wise normalization and pruning yield dense advantages and reduce cost.
@@ 3.3 Reward Fusion and Depth-wise Advantage Estimation
where $s_{\ell}$ is the path log-probability under the behavior policy and $\beta$ controls the strength of weighting. When $\beta=0$, this reduces to simple averaging; larger $\beta$ emphasizes high-likelihood paths and reduces the influence of noisy, low-probability leaves, in all our experiments we set $\beta=1.0$.
@@ 3.4 Pruning Strategies
While branch rollouts improve efficiency and provide dense process rewards, an excessive number of branches may induce exponential growth in trajectory count, leading to prohibitive backpropagation cost. Recent work (Li et al., 2025a) has revealed that standard GRPO training contains substantial redundancy and pruning potential. Motivated by this, we introduce two complementary pruning strategies in the context of BranchGRPO: width pruning, which reduces the number of leaves used for gradient updates, and depth pruning, which skips unnecessary denoising layers.
@@
After computing rewards and advantages for all leaves $\mathcal{L}$, we optionally select a subset of them for gradient updates. We study two simple modes. The first is Parent-Top1, which retains the higher-reward child of each parent in the last branching step. This strategy halves gradient cost while ensuring coverage of all parents, yielding stable but slightly less diverse updates. The second is Extreme selection, which preserves both the globally best and worst $b$ leaves by reward score. By explicitly maintaining strong positive and negative signals, this mode injects more exploratory feedback but introduces higher variance.
@@
Notably, recent work MixGRPO (Li et al., 2025a) achieves strong performance by hybrid ODE–SDE, effectively acting as a form of depth pruning. This motivates our explicit exploration of depth pruning within BranchGRPO.
@@ 4.1 Experiment Setup
We evaluate BranchGRPO on HPDv2.1 (Wu et al., 2023) (103k training and 400 balanced test prompts). The backbone is FLUX.1-Dev, with rollouts augmented by branching and process-dense rewards; we also study a pruned variant (BranchGRPO-Pruning). Baselines include DanceGRPO and MixGRPO under identical settings. We report efficiency (NFE, iteration time, GPU-hours) and quality (HPS-v2.1, PickScore (Kirstain et al., 2023), ImageReward (Xu et al., 2023)).
@@ 4.2 Implementation Details
We set the tree depth to $d=20$ and the branch factor to $b=2$, yielding $16$ leaves per rollout before pruning. The branching steps $\mathcal{B}$ use three presets: Dense $(0,3,6,9)$ as the default, Mixed $(0,5,15,20)$, and Sparse $(0,10,13,19)$. The branch correlation sweeps $s\in\{0,1,2,4,8\}$. Training runs for 300 optimizer steps with gradient accumulation $g=12$ and per-GPU batch size =2=2, on $16\times$ NVIDIA H20z GPUs. Optimization uses AdamW (learning rate $1\times 10^{-5}$, weight decay $1\times 10^{-4}$) with bf16 precision and EMA weights stored on CPU. All GRPO-related hyperparameters are kept identical across methods, with full details deferred to the supplementary material.
Figure 4: Qualitative comparison of generations from Flux, DanceGRPO, and our BranchGRPO.
@@ 4.3 Main Results
Table 1 compares efficiency and alignment quality across methods, while Figure 1 shows reward dynamics during training. BranchGRPO achieves consistently better alignment than DanceGRPO while using substantially fewer function evaluations and shorter iteration times. MixGRPO improves efficiency but at the cost of weaker alignment, whereas BranchGRPO provides a more favorable balance, and pruning variants further reduce computation without hurting performance.
The reward curves in Figure 1 also highlight the stability of our approach. DanceGRPO with reduced timestep fractions suffers from strong reward fluctuations, while the full-timestep variant converges more smoothly but at much higher cost. BranchGRPO combines the best of both, offering faster reward growth and more stable trajectories. Figure 4 further illustrates qualitative improvements: compared with Flux and DanceGRPO, our method produces more faithful, detailed, and coherent generations across diverse prompts.
Taken together, these results demonstrate that BranchGRPO is both more efficient and more stable than existing GRPO variants, yielding superior preference alignment under significantly reduced computational overhead.
@@ 4.4 Ablation Studies
We next vary the positions of split timesteps (Figure 5(b)). Early splits such as $(0,3,6,9)$ promote faster reward increase in the early stage, whereas later splits like $(9,12,15,18)$ delay exploration and yield lower rewards. Intermediate schedules (e.g., $(3,6,9,12)$) balance efficiency and reward quality, suggesting that early splits are generally more effective for exploration.
@@
Figure 5: Ablation on branching hyperparameters without pruning. (a) Moderate correlation (s=4s{=}4) best balances exploration and stability. (b) Early split schedules yield higher rewards than late splits. (c) Denser early branching accelerates reward growth while converging to similar final levels.
@@ 4.5 Video Generation Results
Quantitatively, BranchGRPO achieves faster convergence and higher rewards, measured by the Video-Align motion quality metric (Liu et al., 2025b), confirming the benefit of branching rollouts for video generation.
@@ 5 Conclusion
Appendix A Appendix
This appendix provides additional theoretical proofs, implementation details, and experimental results that complement the main paper. The contents are organized as follows:
  • Section B: Hyperparameter settings used in all experiments.
  • Section C: Theoretical analysis, including proofs for branch noise construction, reward fusion, and variance reduction.
  • Section D: Additional experiments, including ablations, more text-to-image and image-to-video results.
  • Failure Cases: Qualitative examples where BranchGRPO fails, highlighting current limitations.
Appendix B Hyperparameter Settings
Table 2 summarizes the detailed hyperparameter configuration used in our experiments. All hyperparameters are kept identical across all methods, including DanceGRPO and MixGRPO, to ensure a fair comparison.
Table 2: Hyperparameter settings used in all experiments.
Parameter Value Parameter Value
Random seed 42 Learning rate $1\times 10^{-5}$
Train batch size 2 Weight decay $1\times 10^{-4}$
SP size 1 Mixed precision bfloat16
SP batch size 2 Grad. checkpointing Enabled
Dataloader workers 4 Max grad norm 0.01
Grad. accum. steps 12 Warmup steps 0
Checkpoint steps 40 Use TF32 Yes
Resolution $720\times 720$ Sampling steps 16
Eta 0.3 Sampler seed 1223627
Num. generations 12 Shift (branch offset) 3
Use group reward Yes Ignore last step Yes
Clip range $1\times 10^{-4}$ Adv. clip max 5.0
Use EMA Yes EMA decay 0.995
Init same noise Yes
Appendix C Theoretical Analysis (Detailed Proofs)
C.1 Branch Noise Construction and Boundary Distribution Preservation
Consider the reverse SDE discretized by Euler–Maruyama:
$z_{i+1}=\mu_{\theta}(z_{i},t_{i})+g_{i}\,\eta_{i},\qquad\eta_{i}\sim\mathcal{N}(0,I),\ \ g_{i}:=g(t_{i})\sqrt{h_{i}}.$ (6)
At a split step $i\in B$, we construct $K$ branch noises as
$\xi_{b}=\frac{\xi_{0}+s\,\delta\xi_{b}}{\sqrt{1+s^{2}}},\qquad\xi_{0},\delta\xi_{b}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(0,I),$ (7)
so that $\xi_{b}\sim\mathcal{N}(0,I)$ for each $b$, and $\mathrm{Cov}(\xi_{b},\xi_{b^{\prime}})=\tfrac{1}{1+s^{2}}I$ for $b\neq b^{\prime}$.
Each child branch then updates as
$z^{(b)}_{i+1}=\mu_{\theta}(z_{i},t_{i})+g_{i}\,\xi_{b}.$ (8)
Lemma 1 (Single-step marginal preservation).
For any fixed parent $z_{i}$, we have
$z^{(b)}_{i+1}\ \stackrel{{\scriptstyle d}}{{=}}\ \mu_{\theta}(z_{i},t_{i})+g_{i}\eta_{i},\qquad\eta_{i}\sim\mathcal{N}(0,I).$
Lemma 2 (Leaf marginal preservation).
Conditioned on prefix $(z_{0},\eta_{0},\dots,\eta_{i-1})$, each branch $z^{(b)}_{N}$ generated by the split rule has the same distribution as a baseline SDE sample $z_{N}$.
Theorem 1 (Boundary distribution invariance).
Under the branching construction above, for any set of split steps $B$, each leaf $z^{(b)}_{N}$ has the same marginal law as a baseline SDE rollout. Hence, branching does not alter the final generator distribution.
C.2 Reward Fusion: Unbiasedness and Variance Reduction
Let $L(n)$ be the leaf set of a node $n$. Each leaf has reward $r_{\ell}=r(z_{N}^{(\ell)})$. Define the conditional expected return
$V(n):=\mathbb{E}[\,r(z_{N})\mid n\,].$
Uniform fusion.
$\bar{r}(n)=\frac{1}{|L(n)|}\sum_{\ell\in L(n)}r_{\ell}.$ (9)
Then
$\mathbb{E}[\bar{r}(n)\mid n]=V(n),\qquad\mathrm{Var}(\bar{r}(n)\mid n)=\frac{\sigma^{2}(n)}{|L(n)|},$
where $\sigma^{2}(n)=\mathrm{Var}(r_{\ell}\mid n)$.
Path-probability weighted fusion.
If leaves are drawn from proposal $q(\ell\mid n)$ with weights
$w_{\ell}=\frac{p_{\text{beh}}(\ell\mid n)}{q(\ell\mid n)},$
then the IS estimator
$\hat{r}_{\mathrm{IS}}(n)=\frac{1}{|L(n)|}\sum_{\ell\in L(n)}w_{\ell}r_{\ell}$
is unbiased: $\mathbb{E}[\hat{r}_{\mathrm{IS}}(n)\mid n]=V(n)$. The self-normalized IS estimator
$\hat{r}_{\mathrm{SNIS}}(n)=\frac{\sum_{\ell}w_{\ell}r_{\ell}}{\sum_{\ell}w_{\ell}}$
is consistent with variance $O(1/\mathrm{ESS})$, where
$\mathrm{ESS}=\frac{(\sum_{\ell}w_{\ell})^{2}}{\sum_{\ell}w_{\ell}^{2}}.$
C.3 Depth-wise Baseline (Control Variates)
For $K$ siblings at depth $i$, let fused returns $\bar{r}^{(b)}$, and group mean $\bar{r}_{i}=\tfrac{1}{K}\sum_{b}\bar{r}^{(b)}$. Define
$A_{i}^{(b)}=\bar{r}^{(b)}-\bar{r}_{i},\qquad\sum_{b}A_{i}^{(b)}=0.$ (10)
Let $g_{i}^{(b)}(\theta)=\nabla_{\theta}\log p_{\theta}(\text{branch }b\text{ at depth }i)$. Then
$\widehat{\nabla J}_{\mathrm{group}}=\frac{1}{K}\sum_{b=1}^{K}A_{i}^{(b)}\,g_{i}^{(b)}$
is an unbiased gradient estimator with strictly smaller variance than $\widehat{\nabla J}_{\mathrm{single}}=\frac{1}{K}\sum_{b}\bar{r}^{(b)}g_{i}^{(b)}$, unless $\mathrm{Cov}(\bar{r}^{(b)},g_{i}^{(b)})=0$.
C.4 Continuous Rewards and Concentration
Assume $r$ is $L$-Lipschitz and the SDE flow $\Psi_{i+1\to N}$ is $K_{i}$-Lipschitz. Then
$|r(z_{N}^{(b)})-r(z_{N}^{(b^{\prime})})|\leq LK_{i}g_{i}\|\xi_{b}-\xi_{b^{\prime}}\|.$
Thus $r(z_{N}^{(b)})$ is sub-Gaussian with parameter $\mathcal{O}(L^{2}K_{i}^{2}g_{i}^{2})$. Averaging over $|L(n)|$ leaves gives
$\Pr\big{(}|\bar{r}(n)-V(n)|\geq\varepsilon\mid n\big{)}\leq 2\exp\left(-c\cdot|L(n)|\,\varepsilon^{2}/(L^{2}K_{i}^{2}g_{i}^{2})\right).$
Appendix D Additional Experiments
D.1 More quantitative result
Table 3 reports more results.
Table 3: Ablation study of BranchGRPO under different design choices. Best and second-best per column are in bold and underline. All results are obtained under the same training setup as in Section 4.3.
[…]
Branch Density
$(0,3,6,9)$ 13.68 13.68 493 0.363 0.229 1.233 0.374
$(0,4,8,12)$ 11.56 11.56 416 0.359 0.229 1.238 0.374
$(0,5,10,15)$ 9.44 9.44 340 0.354 0.228 1.220 0.366
Reward Fusion
Uniform Fusion 13.68 13.68 493 0.361 0.229 1.218 0.368
Path-Weighted Fusion 13.68 13.68 493 0.363 0.229 1.233 0.374
Pruning Strategy
Width Pruning 13.68 8.625 314 0.364 0.230 1.300 0.374
Depth Pruning 13.68 8.625 314 0.369 0.231 1.319 0.381
D.2 More Text2Image Results
Figure 8:
Figure 9:
Figure 10:
Figure 11:
Figure 12:
D.3 More Image2Video Results
(a) Case 1
(b) Case 2
(c) Case 3
(a) Case 4
(b) Case 5
D.4 Failure Cases
Figure 15: Failure case
Figure 16: Failure case